in triangle DEF, m angle E is three times m angle D, and m angle F is 9 less than m angle E. what is the measure of each angle?

Answer by ikleyn(52748)

.

The sum of the angles is 180 degrees


    D +  E +  F     = 180

    D + 3D + (3D-9) = 180

    7D              = 189

     D              = 189/7 = 27.


From this point, complete the solution on your own.

Standard trivial routine problem.